OpenStudy (anonymous):

use cylindrical coordinates to find the volume of the region in the first octant bounded by cylinder x^2+y^2=4 and the plane z=1-2x+3y

ganeshie8 (ganeshie8):

\(\large \iiint \limits_R dz ~r dr d\theta \)

OpenStudy (anonymous):

so change to polar, huh..

ganeshie8 (ganeshie8):

\(\large \int \limits_0^{\frac{\pi}{2}} ~ \int \limits_0^{2}~~\int \limits_0^{ 1-2r\cos\theta + 3r\sin\theta } dz ~r dr d\theta \)

ganeshie8 (ganeshie8):

evaluate

OpenStudy (anonymous):

ok, thanks! I just needed help setting up! so it's easier to use polar when you see something like \[x^2 + y^2 \]. but what does "use cylindrical coordinates" mean?

ganeshie8 (ganeshie8):

polar = 2D cylinderical = 3D

OpenStudy (anonymous):

thank you for the help!

ganeshie8 (ganeshie8):

polar : \(\large (r, \theta)\) cylindrical : \(\large (r, \theta, z)\)

ganeshie8 (ganeshie8):

you may simply setup the double integral for volume, if you're not familiar wid triple integrals...

ganeshie8 (ganeshie8):

\( \mathbb{ \int \limits_0^{\frac{\pi}{2}} ~ \int \limits_0^{2}~~\int \limits_0^{ 1-2r\cos\theta + 3r\sin\theta } dz ~r dr d\theta = \int \limits_0^{\frac{\pi}{2}} ~ \int \limits_0^{2}~~ (1-2r\cos\theta + 3r\sin\theta ) ~r dr d\theta }\)

ganeshie8 (ganeshie8):

right side is just a double integral

OpenStudy (anonymous):

would that give me the same answer? because I am more familiar with double integrals.

ganeshie8 (ganeshie8):

yup ! ive just evaluated the inner integral in triple integral :)

ganeshie8 (ganeshie8):

\(\mathbb{ \int \limits_0^{\frac{\pi}{2}} ~ \int \limits_0^{2}~~\color{red}{\int \limits_0^{ 1-2r\cos\theta + 3r\sin\theta } dz} ~r dr d\theta = \int \limits_0^{\frac{\pi}{2}} ~ \int \limits_0^{2}~~\color{red}{ (1-2r\cos\theta + 3r\sin\theta )} ~r dr d\theta }\)

ganeshie8 (ganeshie8):

the integrand u get when u do double integral is simply the bound for inner integral in triple integral